# UnlearnCanvas: A Stylized Image Dataset to Benchmark Machine Unlearning for Diffusion Models

### v2

@@ C.2 Broader Use Cases of UnlearnCanvas
Bias mitigation. Bias mitigation in DMs, which are now gaining popularity, can also benefit from UnlearnCanvas. Its hierarchical and balanced architecture enables the deliberate introduction of artificial biases by selectively omitting data from specific groups. For instance, by predominantly excluding images from styles other than the ‘Van Gogh Style’ within the ‘Dogs’ class, DMs fine-tuned on this dataset will inherently exhibit a tendency to generate images of dogs in the Van Gogh style, particularly when the style is not explicitly specified in the prompt. This approach not only allows for the manipulation and quantification of biases but also paves the way for UnlearnCanvas to become a standardized benchmark for bias mitigation, similar to the role of MU for DMs.
Vision in-context learning (V-ICL). V-ICL (Wang et al., 2023b, a; Bai et al., 2023; Wang et al., 2024) is another domain where UnlearnCanvas can be effectively applied. The field of V-ICL is in urgent need of robust, comprehensive methods for the fair assessment of existing models. In this context, the image pairs from UnlearnCanvas are ideally suited for evaluating various tasks such as style transfer, image inpainting, and image segmentation, offering a rich resource for nuanced and quantitative analyses.

### v4

@@ F.2 Other Possible Applications of UnlearnCanvas
Bias mitigation. Bias mitigation in DMs, which are now gaining popularity, can also benefit from UnlearnCanvas. Its hierarchical and balanced architecture enables the deliberate introduction of artificial biases by selectively omitting data from specific groups. For instance, by predominantly excluding images from styles other than the ‘Van Gogh Style’ within the ‘Dogs’ class, DMs finetuned on this dataset will inherently exhibit a tendency to generate images of dogs in the Van Gogh style, particularly when the style is not explicitly specified in the prompt. This approach not only allows for the manipulation and quantification of biases but also paves the way for UnlearnCanvas to become a standardized benchmark for bias mitigation, similar to the role of MU for DMs.
Vision in-context learning (V-ICL). V-ICL [93, 94, 95, 96] is another domain where UnlearnCanvas can be effectively applied. The field of V-ICL is in urgent need of robust, comprehensive methods for the fair assessment of existing models. In this context, the image pairs from UnlearnCanvas are ideally suited for evaluating various tasks such as style transfer, image inpainting, and image segmentation, offering a rich resource for nuanced and quantitative analyses.
G Impact Statement
$\bullet$ Avoiding Copyright Issues. There is an urgent need for the generative model providers to scrub the influence of certain data on an already-trained model. In January 2023, a notable lawsuit targeted two leading AI art generators, Stable Diffusion [1] and Midjourney [2], for alleged copyright infringement. Concurrently, incidents with the recently released Midjourney V6 [2] also highlighted a visual plagiarism issue on famous film scenes. These instances illuminate the broad copyright challenges inherent in the way of training data collection method of those foundation generative models’ training datasets. MU methods can be used as an effective method to remove the influence of the private data and avoid unnecessary retraining.
$\bullet$ Mitigating biases and stereotypes. Generative AI systems are known to have tendencies towards bias, stereotypes, and reductionism, when it comes to gender, race and national identities [17]. For example, a recent study on the images generated with Midjourney revealed, that images associated with higher-paying job titles featured people with lighter skin tones, and that results for most professional roles were male-dominated [16]. MU is known to be effective in eliminating biases rooted in the training data. Moreover, UnlearnCanvas offers a flexible framework to benchmark MU techniques against bias removal, allowing for the creation and quantitative control of biases across different object classes for comprehensive bias removal studies.